OpenStudy (anonymous):

What is the percent of decrease on an item that went from $25 to $20?

OpenStudy (anonymous):

So here we go: \[\frac{Decrease}{New \; Amount} \times 100\]

OpenStudy (anonymous):

I means if we have to find the decrease percentage between x to y : Then: \[\implies \frac{x - y}{y} \times 100\]

OpenStudy (anonymous):

*mean.

OpenStudy (anonymous):

(x + 2) + 10

OpenStudy (anonymous):

What??

OpenStudy (anonymous):

Here x = 25 and y = 20, so use the formula that I have given above..